### v2

# Mercury: A Computational Efficiency Benchmark for Code Large Language Models

Abstract
Amidst the recent strides in evaluating Large Language Models for Code (Code-LLMs), existing benchmarks have mainly focused on functional correctness, overlooking the importance of computational efficiency. To fill the gap, we present Mercury, the first computational efficiency benchmark for Code-LLMs. It comprises 1,889 Python tasks, each with adequate solutions to support a runtime distribution. Based on the distribution, we introduce a new metric $Beyond$, which computes a runtime-percentile-weighted $Pass$ score to reflect functional correctness and computational efficiency simultaneously. On Mercury, leading Code-LLMs can achieve 67% on $Pass$, while less then 50% on $Beyond$. Given that an ideal $Beyond$ score would be aligned with the $Pass$ score, it indicates that while Code-LLMs exhibit impressive capabilities in generating functionally correct code, there remains a notable gap in their efficiency. Finally, our empirical experiments reveal that Direct Preference Optimization (DPO) serves as a robust baseline for enhancing computational efficiency compared with Supervised Fine Tuning (SFT), which paves a promising avenue for future exploration of efficient code generation. [^1] Our code and data are available on GitHub: https://github.com/Elfsong/Mercury..
1 Introduction
The domain of code generation, which aims to empower computers to autonomously generate code based on natural language task instructions (NL2Code), constitutes a longstanding and fascinating research objective [42]. The recent emergence of Large Language Models (LLMs) has spurred the development of a new wave of code generation models [40, 31, 35, 11, 27, 4], which leverage the impressive language understanding and generative capabilities of LLMs to drive forward the ambitious goal of producing high-quality and machine-generated code.
In parallel with the model development, recent code generation benchmarks mainly focus on evaluating functional correctness in the fuzz testing style [25]. This testing approach assesses the outcome congruence between the generated and anticipated solutions by executing bespoken test cases. For instance, HumanEval [8] and MBPP [3] collected a small but high-quality set of handcrafted tasks, and EvalPlus [24] further consolidates these two above benchmarks by augmenting the test scope. On the contrary, APPS [14] widely gathered over 5,000 public coding tasks from online platforms. Despite these strides, there is a discernible oversight in current code generation benchmarks concerning the computational efficiency evaluation, although that is critical in real-world software development [43, 44]. Moreover, for most tasks in the above benchmarks, handcrafting diverse solutions and comprehensive test cases to cover all possible scenarios is a time-consuming endeavor [24]. In light of these findings, we highlight several vital limitations inherent in the existing code generation benchmarks:
Figure 1: While both LLM-generated codes successfully follow the task instruction, the right snippet notably excels in computational efficiency, executing in a mere 121 ms compared to the 5,714 ms consumed by the left snippet. On the premise of ensuring functional correctness, optimizing computing efficiency is a ubiquitous objective in software development.
  1. 1.
    Absence of Computational Efficiency Evaluation. Existing code generation benchmarks focus on assessing functional correctness while overlooking the evaluation of computational efficiency [8, 3, 14]. As illustrated in Figure 1, despite both code snippets can handle the sorting task functionally, the right efficient solution (121 ms) is nearly 50 times faster than the left inefficient solution (5,714 ms). This striking runtime differentiation underscores the necessity of incorporating computational efficiency assessments within code generation benchmarks to ensure that synthesized code is not only correct but also efficient.
  2. 2.
    Insufficient Test Case Coverage. Most code generation benchmarks obtain a limited set of handcrafted test cases for assessing functional correctness, potentially overlooking edge cases and nuanced code behaviors [8, 3]. For example, Figure 3 displays that HumanEval #55 contains only 3 test cases, testing up to the 12th Fibonacci number [8]. Its given canonical solution will quickly reach the recursion depth limitation when computing a larger Fibonacci number (the recursion limitation depends on the environment). Therefore, notwithstanding the generated code satisfies all test cases, such success does not necessarily equate to assurance of functional correctness and much less to computational efficiency.
  3. 3.
    Lack of Task Diversity. Another noticeable deficit of existing code generation benchmarks is the insufficient diversity and complexity in their constituent tasks [8, 3, 25]. Since most benchmarks only consist of elementary-level problems, recent Code-LLMs can effortlessly tackle most tasks regardless of their actual capacities [44]. This flaw results in these benchmarks failing to pose a substantial challenge to Code-LLMs and truly reflect their underlying potential.
Computational Efficiency.
Computational efficiency refers to the performance measure of time and space complexity to accomplish a specific task. Efficient code can improve user experience, save energy, and make applications more sustainable and cost-effective. Compared with the scalable memory space, execution time is the performance bottleneck of most programs. Consequently, this work focuses on the time dimension of computational efficiency.
Figure 2: An overview of Mercury dataset. Each Mercury task has a task description, a test case generator, a prompt & entry point, and corresponding solutions. To evaluate computational efficiency, we introduce the $Beyond$ metric, which signifies the runtime percentile of the LLM-generated code on the runtime distribution supported by corresponding solutions. In this example, the LLM-generated code executes in 521 ms, outpacing 86.18% of collected solutions on the runtime distribution. Consequently, the $Beyond$ metric in this case is 86.18%.
Our Benchmark.
In this work, we introduce Mercury, a novel code generation benchmark designed to assess computational efficiency. As depicted in Figure 2, Mercury comprises 1,889 Python programming tasks with three difficulty stratification, which is divided into two datasets for model fine-tuning and evaluation separately. For each task, we assign a test case generator to remedy the shortfall of test case coverage. In measuring computational efficiency, the primary challenge stems from normalizing the absolute runtime across tasks that have diverse runtime ranges. Thus we collect and locally execute numerous solutions for each task to form a runtime distribution and leverage the runtime percentile of LLM-generated code on the distribution instead of the absolute runtime to evaluate computational efficiency. Furthermore, to mitigate performance discrepancies attributed to irrelevant processes and diverse hardware configurations, we set up an isolated sandbox environment for task execution to establish local runtime distributions.
Contribution.
Our work aimed to fill the computational efficiency evaluation gap in code generation benchmarks with the following key contributions:
  • Dataset. We introduce a novel code generation dataset Mercury designed to assess and improve computational efficiency in Section 2, accompanied by an extensible open-source data collection framework for enriching Mercury with more tasks and languages.
  • Metric. We propose the first efficiency-focused code generation metric $Beyond$ and establish a benchmark to evaluate leading Code-LLMs using this metric in Section 3.
  • Baselines. In Section 4, we detail our extensive analysis of baselines to enhance computational efficiency while maintaining functional correctness. Experiment results reveal that despite Code-LLMs excelling in functional correctness, there is still considerable potential to elevate efficiency.
Table 1: A comparison of Mercury to existing NL2Code benchmarks. Mercury distinguishes itself by including a set of distilled high-quality solutions and a dedicated test case generator for each task.
Benchmarks Tasks Sources Cases Solutions Difficulty
HumanEval 164 Crowd Source 8.08 1 1
MBPP 257 Crowd Source 3.01 1 1
APPS 5,000 Online 21.2 23.4 3
Mercury 256 Online + Filters $\infty$ 18.4 3
Figure 3: An example of insufficient test cases. Even though the code passed all test cases in the dashed-line box, it remains vulnerable to timeout or stack overflow when subjected to a larger input.
2 Mercury Datasets
We initiate this work by collecting public programming tasks on Leetcode [^2] https://leetcode.com/problemset/algorithms/. Subjecting these questions to a series of filters, we distilled them down to 1,889 high-quality tasks. A difficulty-balanced subset of 256 tasks was randomly selected to form the Mercury-eval benchmark, which obtains an average of 18.4 solutions for each problem. The remaining tasks have been designated as the Mercury-train dataset for baseline training (detailed data distribution is listed in Appendix Table 6). To enhance clarity within this paper, we employ Mercury to denote Mercury-eval unless otherwise specified.
Data Schema.
As illustrated in Figure 2, Mercury offers a unified data schema to streamline the evaluation procedure and bolster further development endeavors. The data scheme encompasses these principal components: (1) Task Description contains the task instruction interpreted into a plain text format, along with illustrative examples and constraints of inputs and outputs. (2) Test Case Generator refers to a Python code snippet designed to automatically produce a comprehensive set of test cases in accordance with the specifications laid out in the task description. (3) Solutions are sampled from Leetcode historical submissions. Each solution within Mercury has undergone rigorous testing, and Locality-Sensitive Hashing [16] is employed to prevent the inclusion of any identical solutions. (4) Prompts and Entry Points where prompts act as the initiating prefixes for LLM code generation and entry points denote the start point for code execution. We delineate the definition of Mercury fields in the Appendix Table 5.
Task Filters.
Mercury tasks originate from public programming problems on Leetcode. To assure the quality and uniformity of the dataset, we distilled gathered tasks based on the following conditions:
Number of Solutions. To establish a solution runtime distribution for each task, we filtered out tasks having less than two associated solutions. After excluding these tasks, Mercury tasks possess an average of 18.4 unique solutions.
Restricted Data Structure. Above the inherent Python data types, Mercury also incorporates two custom data types: Binary Tree and Linked List (the specific structure definitions can be found in Figure 4), which increases Mercury’s diversity and escalates its difficulty level. Tasks that contain other data structures will be removed.
Unique Outputs. Certain Leetcode tasks may permit non-unique answers. For example, a result list can be returned in any order. Evaluating all possible answers can drastically complicate the test case verification process. To eliminate this problem, we harness the corresponding test case generator to generate $N$ test cases $T_{i}=\langle Input_{i},Output_{i}\rangle\ s.t.\ i\in\{0,1,\cdots,N\}$ and execute $T$ on different solutions $S_{m}\ s.t.\ m\in\{0,1,\cdots,M\}$ to observe if all outputs $Output_{i}=S_{m}(Input_{i})$ remain identical. Any tasks that potentially yield non-unique answers were subsequently excluded.
Task Difficulty.
Most existing NL2Code benchmarks predominantly comprise simplistic tasks, leading to a situation where LLMs of varied capabilities address most tasks effortlessly and yield indistinguishable high scores [44, 17]. To alleviate this issue, Mercury inherits the difficulty categorization from Leetcode, i.e., Easy, Medium, and Hard. The stratification aims to probe the upper bounds of Code-LLM capabilities, delivering a more evident distinction between various Code-LLMs.
Test Case Generator.
Manual creation of test cases can be a laborious process. To gather sufficient test cases to conduct an exhaustive assessment, we assign a test case generator for each task, which can produce a full range of test cases to thoroughly evaluate the functional correctness and efficiency of given solutions. Specifically, We reformulate the task description as plain text (see $pretty\_content$ in Table 5) and then feed it into GPT-4 [31] to generate an initial test case generator snippet. To confirm the effectiveness of the initial generator, we subsequently create a batch of test cases (24 cases in our setting) by the generator and submit them to the Leetcode online judge system. Should any of the generated test cases not pass the Leetcode validation, we manually revise the generator until all generated cases can be successfully validated.
3 Computational Efficiency Metric
In the domain of software development, computational efficiency can be defined as the absolute code runtime for executing a given test case set [7]. Nonetheless, a primary obstacle in benchmarking computational efficiency is normalizing runtime measurements across disparate environments. For instance, a sub-optimal solution might have a faster absolute runtime on high-performance hardware than an optimal solution on low-performance hardware. Moreover, different operation systems and code interpreters may also fluctuate the code runtime. Therefore, absolute runtime fails as a consistent and reliable computational efficiency benchmark metric. To address this issue, an intuitive approach involves modeling a devoted runtime distribution for each task and calculating the average runtime percentiles of LLM solution samples over the runtime distribution. With this idea in mind, we proposed a normalized computational efficiency metric $Beyond$:
$p_{k}^{n}=\frac{max(R^{n})-clip(r_{k}^{n},min(R^{n}),max(R^{n}))}{max(R^{n})-min(R^{n})},\qquad Beyond=\frac{\sum\nolimits_{N,K}^{n=0,k=0}p_{k}^{n}}{N\cdot K}.$ (1)
Where $N$ is the total number of tasks, and $K$ denotes the size of LLM solution samples. For a specific task $n\in N$, $R^{n}$ is the runtime array corresponding to the collected solutions, and $r_{k}^{n}$ denotes the runtime for the $k$-th LLM solution. Runtime is defined as the period from the solution instantiation to the evaluation across all test cases, culminating with a successful termination (More engineering details can be found in Appendix Figure 5). Since any case failure of the $k$-th solution results in $r_{k}^{n}\rightarrow+\infty$ and then $p_{k}^{n}=0$, $Beyond$ reflects functional correctness as well.
Untrusted Code Execution.
Since most Code-LLMs are trained on an extensive code corpus from unverified sources, there is an intrinsic risk that these models may produce malicious code when driven by specific meticulous prompts [8]. The direct execution of synthesized code raises significant security concerns. To alleviate the risk of running untrusted code, we engage a robust sandbox to execute code in an isolated environment. Sandbox details are deliberated in Appendix A.3
Hardware-agnostic Evaluation.
To ensure fair comparison across diverse hardware configurations, we run each task $n$ with corresponding test cases locally and aggregate their runtimes into the runtime array $R^{n}$. Appendix Figure 8 illustrates the $Beyond$ score of two LLMs (‘deepseek-coder-33b’ and ‘deepseek-coder-6.7b’) over three distinct hardware specifications: the micro-tier (0.25 CPU cores), the small-tier (0.5 CPU cores), and the standard-tier (1 CPU core). The results demonstrate that Beyond remains consistent over different hardware configurations.
4 Experiments
In this section, we present a series of baseline experiments to improve computational efficiency by training on Mercury-train dataset and assessing on the Mercury-eval dataset. Our empirical study encompasses 10 open-source LLMs a broad parameter spectrum from 1.3 to 34 billion. For each LLM, we compare the performance of the original model and two optimization strategies: Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), for their potential to optimize LLM generating functionally correct and computationally efficient code. Finally, we analyzed the underlying factors contributing to the failure of LLMs on the Mercury-eval dataset.
4.1 Baselines
Supervised Fine-Tunning. Within the SFT [6] method, an LLM undergoes additional training on a small dataset, which aims to specialize the LLM to perform better on certain tasks correlated to the training dataset. To optimize the computational efficiency performance of Code-LLMs, the most intuitive strategy is to fine-tune the Code-LLM using optimal runtime solutions. In our experimental setup, we apply the unified prompt template for each Code-LLM to ensure a fair comparison. The “pretty_content” attribute fills the <task_content> placeholder, the “prompt” attribute fills the <code_starter> placeholder, and the <code_completion> placeholder is completed with the fastest solutions. To steer Code-LLMs towards generating the intended code completion, we prepend a one-shot example to the prompt template. Appendix Figure 9 presents the prompt template.
Direct Preference Optimization. Although SFT exemplifies a straightforward approach, it is susceptible to the pitfall of catastrophic forgetting [19]. To enable LLMs to align with human preferences while preserving their functional capabilities, existing methodologies employ reinforcement learning with human preference feedback (RLHF). However, RLHF introduces additional complexities and potential instabilities, necessitating significant computing resources and additional reward model training [45, 5, 37]. DPO [33] bypasses these challenges by explicitly mapping reward functions and the optimal objective. This connection demonstrates that maximizing rewards under specific constraints can be effectively addressed through a singular training phase based on data reflecting human preferences. Consequently, DPO provides a stable, efficient, and computation-friendly alternative, removing the necessity for separate reward modeling. The DPO dataset construction and training procedure are elaborated in Appendix Section A.4.
4.2 Functional Correctness Benchmarks
HumanEval assesses the functional correctness of synthesized code derived from docstrings. It contains 164 distinct Python tasks that cover several programming areas, such as language comprehension, algorithm development, and simple mathematics [8]. MBPP has a sanitized collection of 257 entry-level Python programming problems. Each problem in this dataset consists of three components: a task description, an associated code solution, and three automated test cases to validate the code functionality [3]. Both HumanEval and MBPP harness the metric $Pass$ to measure the Code-LLMs’ functional correctness, where a task is considered solved if it passes all test cases, and the total fraction of solved passed is reported [20].
4.3 Experimental Setups
Configuration.
[…]
Training Data.
We use Mercury-train for model training. As for the SFT process, we nominate the fastest solution as the supervised label, then format the training data as $\langle pretty\_content,prompt,solution\_optimal\rangle$. Regarding the DPO procedure, we select the top 5 pairs of solutions that exhibit the most significant discrepancy in runtime. The training date format is $\langle pretty\_content,prompt,solution\_fast,solution\_slow\rangle$.
4.4 Empirical Results
Functional correctness is the prerequisite for evaluating computational efficiency for the code generation task. Our primary objective is to enhance the computational efficiency without compromising the functional correctness. To this end, we first introduce the existing metric $Pass$ to gauge the functional correctness [8] and then leverage $Beyond$ to provide a holistic evaluation, encompassing both computational efficiency and functional correctness. Finally, we measure the $Gap$ between $Beyond$ and $Pass$ to indicate the baseline effectiveness. These experiments aim to scrutinize the innate capabilities of cutting-edge Code-LLMs and their potential after baseline fine-tuning. Therefore, extensive parameter optimization and prompt engineering were not pursued. To deliver a comprehensive evaluation, we have further integrated the HumanEval and MBPP benchmarks as supplementary measures for appraising functional correctness [8, 3].
Functional Correctness.
Table 2 lists functional correctness benchmarks over various models, showing that larger models tend to provide better functional correctness. Except for the smallest model “deepseek-coder-1.3b-base”, DPO consistently enhances the overall $Pass$ scores across most Code-LLMs, while SFT diminishes functional correctness on the largest two Code-LLMs. These findings suggest that smaller models may struggle to integrate new knowledge while preserving their original functionality, and SFT may induce catastrophic forgetting in larger models in the pursuit of heightened computational efficiency. Moreover, it is observable that the functional correctness metrics of each model consistently diminish as the difficulty level increases, where Mercury difficulty stratification probes the model’s upper capability better than the other two benchmarks.
Table 2: Functional correctness ($Pass$) evaluation results over LLMs. The underlined values denote the top-performed approaches among the original model and the established baselines. The bolded values denote the best performance on each benchmark.
Mercury
Model Name HumanEval MBPP Easy Medium Hard Overall
[…]
starcoder2-3b 31.7 57.4 56.1 52.1 21.6 37.8
[…]
deepseek-coder-6.7b-base 47.6 70.2 69.3 68.9 56.1 61.0
[…]
starcoder2-7b 35.2 54.4 63.6 61.7 29.2 44.3
+ SFT 42.9 57.2 64.8 58.5 31.3 47.5 (+3.2)
[…]
+ SFT 29.5 47.6 58.9 38.5 16.1 31.3 (+1.7)
[…]
starcoder2-15b 46.3 66.2 69.5 65.4 50.3 58.0
+ SFT 51.6 69.2 72.0 68.9 51.7 61.3 (+3.3)
[…]
+ SFT 39.5 59.8 65.5 54.8 19.5 39.5 (-0.1)
[…]
deepseek-coder-33b-base 54.3 73.2 70.9 67.9 62.3 65.0
+ SFT 58.1 74.8 61.8 58.0 47.1 58.7 (-6.3)
[…]
CodeLlama-34b-hf 48.2 65.4 77.7 63.7 32.4 52.4
+ SFT 52.8 68.2 61.8 58.0 26.2 47.5 (-4.9)
[…]
Code Efficiency.
Table 3: Computational efficiency ($Beyond$) evaluation results across three difficulty levels. The bolded value indicates the top performance for each metric, while the underlined values denote the most effective approaches among the original model and the baselines.
Model name Easy Medium Hard Overall Gap
[…]
+ SFT 42.58 38.12 18.67 33.04 (-2.58) 12.74 (+2.89)
[…]
+ DPO 43.70 41.02 12.99 32.42 (-0.99) 16.33 (6.61)
[…]
+ SFT 51.37 52.71 44.28 49.39 (+1.09) 16.55 (0.16)
[…]
+ SFT 39.75 26.89 9.55 25.41 (-2.04) 12.48 (3.21)
[…]
+ SFT 54.16 51.43 38.05 47.82 (0.04) 14.91 (-0.43)
[…]
starcoder2-15b 58.18 52.09 37.34 49.17 12.55
+ SFT 53.54 52.77 37.73 47.92 (-1.25) 16.22 (3.67)
[…]
CodeLlama-13b-hf 57.00 44.25 12.99 38.01 13.79
+ SFT 44.95 39.96 13.55 32.70 (-5.31) 13.78 (-0.01)
[…]
deepseek-coder-33b-base 51.26 48.90 45.43 48.53 18.50
+ SFT 40.33 37.75 36.82 38.32 (-10.21) 17.30 (-1.20)
[…]
CodeLlama-34b-hf 56.28 48.21 22.96 42.40 15.49
+ SFT 45.49 44.96 20.73 36.91 (-5.50) 11.61 (-3.88)
[…]
Regarding the NL2Code task, once functional correctness has been assured, attention naturally pivots to enhancing computational efficiency. As depicted in Table 3, we investigate computational efficiency metrics across a spectrum of Code-LLMs. Experiments demonstrate that DPO yields a stable enhancement in computational efficiency from models exceeding 6.7 billion parameters. In contrast, SFT detracts most $Beyond$ scores from original models, suggesting that the plain SFT may not be a feasible strategy for enhancing computational efficiency. Further analysis compares the $Gap$ between $Beyond$ and $Pass$. Since the ideal $Beyond$ should be aligned with $Pass$, it shows how much the baseline method shrinks the gap between functional correctness and computational efficiency. Our findings indicate that DPO substantially narrows $Gap$ in models larger than 15 billion parameters. However, $Gap$ tends to widen in smaller models under the same configuration. This implies that larger models possess a greater capacity to assimilate the nuanced knowledge to make strides in computational efficiency while retaining their functional correctness.
4.5 Failure Analysis
Table 4: The distribution of failure cases across Code Generation, Code Execution, and Test Case errors. E/M/H indicates Easy/Medium/Hard levels, respectively.
Model Name Code Generation Code Execution Test Case Pass
E M H E M H E M H E M H
deepseek-coder-1.3b-base 82 85 59 17 33 95 74 73 180 267 214 101
+ SFT 106 104 37 16 8 63 59 76 225 259 217 110
+ DPO 90 108 40 17 5 49 63 75 259 270 217 87
starcoder2-3b 107 102 33 35 26 107 51 66 201 247 211 94
+ SFT 97 93 24 29 13 90 47 61 211 267 238 110
+ DPO 79 75 11 30 14 87 56 69 235 275 247 102
deepseek-coder-6.7b-base 107 101 30 16 5 56 12 20 105 305 279 244
+ SFT 105 100 25 17 6 61 14 10 98 304 289 251
+ DPO 87 82 23 12 6 58 15 23 98 326 294 256
starcoder2-7b 107 101 22 21 9 74 32 45 212 280 250 127
+ SFT 105 100 22 18 13 72 32 55 205 285 237 136
+ DPO 90 90 21 10 11 61 11 33 211 329 271 142
CodeLlama-7b-hf 23 28 23 41 69 122 131 139 234 245 169 56
+ SFT 11 9 17 44 72 112 126 168 236 259 156 70
+ DPO 9 10 12 23 56 117 111 154 228 297 185 78
CodeQwen1.5-7B 105 100 18 1 4 44 26 17 157 308 284 216
+ SFT 105 101 16 4 3 35 19 26 168 312 275 216
+ DPO 98 96 26 5 8 35 18 30 175 319 271 199
starcoder2-15b 105 100 20 4 7 49 25 33 147 306 265 219
+ SFT 104 100 18 3 3 56 16 23 136 317 279 225
+ DPO 83 64 10 1 1 33 13 41 141 343 299 251
CodeLlama-13b-hf 10 14 28 19 41 99 73 105 228 338 245 80
+ SFT 46 52 32 29 19 111 77 112 207 288 222 85
+ DPO 24 9 24 10 12 100 60 141 185 346 243 126
deepseek-coder-33b-base 105 103 26 11 11 47 12 16 91 312 275 271
+ SFT 69 78 27 27 26 65 72 66 138 272 235 205
+ DPO 56 75 15 9 7 86 28 13 66 347 310 268
CodeLlama-34b-hf 22 35 50 28 55 84 48 57 160 342 258 141
+ SFT 35 97 50 37 19 56 96 54 215 272 235 114
+ DPO 4 12 10 26 76 30 41 40 120 369 277 275
Table 4 provides an error breakdown of where Code-LLMs misstep during the Mercury evaluation:
Generation Errors arise from syntactical issues. The common manifestations include improper indentation, mismatched parentheses, or unexpected truncation. Fine-tuning introduces additional knowledge for Code-LLMs to adapt the Mercury convention, emphasizing standard indentation, concise code, and minimal comments. Therefore, both SFT and DPO generally reduced these errors.
Execution Errors differ from Generation Errors because they occur after the code has been successfully loaded. These errors emerge as exceptions, which could stem from various issues, such as flawed code logic, execution timeouts, memory leakage, or sandbox interruption. We observe that SFT tends to aggravate these errors on most models, whereas DPO mitigates these errors successfully.
Test Case Errors are the most prevalent errors where the code is executed without exceptions but the output fails to align with the expectation.
5 Related Work
NL2Code Generation is the task of generating a computer program that satisfies given specifications. Initial approaches to converting natural language to code relied on rigid methods like probabilistic grammars and domain-specific languages, having limited flexibility and scalability [18, 10]. The advent of statistical models, such as n-grams and Hidden Markov models, attempted to overcome these limitations but struggled with modeling complexity and dependencies [29, 38]. The transformational impact of the Transformer model [39] and its subsequent application to NL2Code [28] led to the development of LLMs like Codex, which significantly improved the task’s feasibility by utilizing extensive unlabelled data sets [8]. Follow-up LLMs such as AlphaCode [23], CodeGen [30], PaLM-Coder [9], and StarCoder [22] continued to advance this research field, exhibiting emergent abilities in coding and debugging that mirrored human programmers.
NL2Code Evaluation currently focuses on gauging the functional correctness of generated code. As a pioneer, CodeBLEU [34] adapts the BLEU [32] metric into code generation. However, given the abstract nature of programming languages, distinct code can express the equivalent semantics, prompting subsequent benchmarks to harness test cases instead of the similarity measurement. For example, HumanEval [8] and MBPP [3] consist of hand-written Python programming tasks and corresponding test cases. On the note of enhancing language inclusiveness, ODEX [41] integrates multiple natural languages, while MBXP [2] extends the benchmarks to cater to a variety of programming languages, promoting polyglot code generation evaluation. Recent benchmarks have also begun to consider more aspects beyond functional correctness. For instance, the benchmark DS-100 [21] dives deeply into the data analysis scenarios, and CodeGen [30] contributes a benchmark for multi-turn code generation. For security-oriented code generation, SecurityEval [36] offers a concentrating benchmark on mining the vulnerability of generated code.
6 Limitations
In this work, we measure computational efficiency under the assumption that the code runtime is uniformly distributed. The simplification allows computational efficiency evaluation via limited solution samples. However, the distribution of code runtime in real-world scenarios is more intricate, which may call for more solution samples to support more precise modeling. Additionally, the presence of data contamination during the model training phase compromises the precision of the Mercury benchmark to reflect the performance of tainted models [17]. To mitigate this issue, we will update our benchmark via our open-sourced data collection framework to import new tasks dynamically, thus laying the groundwork for more detailed investigations in subsequent studies.
7 Conclusion
In this work, we introduced Mercury, the first computational efficiency benchmark for NL2Code evaluation. Unlike prior work that focused on functional correctness, our benchmark highlights the importance of computational efficiency. By crafting dedicated test case generators and sampling ground-truth solutions across different difficulty levels from Leetcode, we have developed a comprehensive and rigorous Code-LLM evaluation frame. We evaluated leading Code-LLMs against benchmarks and found that even though these models are proficient in generating functionally correct code, there is still considerable space for computational efficiency improvement. As Code-LLMs become more widely used, computational efficiency determines factual productivity, where Mercury can gauge the vital metric. As a commitment to ongoing research and to foster further innovation in this area, we have open-sourced the Mercury dataset collection framework, laying the groundwork for future advancements in the field.
Appendix A Appendix
A.1 Definition of Mercury Fields
Table 5: Definitions of the fields within the Mercury dataset.
Field Name Definition
id Task ID.
slug_name Task name.
meta_info The field accommodating the task description and submission statistics.
difficulty The difficulty level of the task.
pretty_content The field introduces the task description, examples, and constraints in pure text.
solutions Samples of solutions extracted from actual past submissions.
prompt The prompt of the solution.
entry_point The nominative entry point of the solution.
generator_code A function to generate test cases.
test_cases A collection of generated test cases.
convert_online A function to format test cases for online evaluation.
convert_offline A function to format test cases for offline evaluation.
evaluate_offline A function designed to evaluate solutions in an offline setting.
A.2 Mercury Data Distribution and Customized Data Structures
Except for all built-in Python data structures, Mercury imports another two structures to enhance the diversity and complexity as shown in Figure 4.
Figure 4: Mercury supports two customized data structures: TreeNode and ListNode.
Table 6: Mercury-eval encompasses 256 tasks, the difficulty of which has been balanced for model evaluation. Mercury-train comprises the remaining 1,633 tasks for model training.
Splits Easy Medium Hard Sum
Mercury-train 446 968 219 1,633
Mercury-eval 88 81 87 256
A.3 Sandbox Details
Time and Memory Limitation.
Each executed code within the sandbox is subject to certain constraints to ensure fair utilization of resources and to prevent any single program from monopolizing the system. Specifically, there are two primary constraints: a time limit and a memory limit. The time limit restricts how long the code can execute before being forcibly terminated, thereby ensuring that no infinite loops or excessively long computations negatively impact the availability of the sandbox. The memory limit caps the amount of RAM that a process can consume. This measure prevents a single program from exhausting the server’s memory resources, which could lead to a denial of service for other users or processes sharing the same infrastructure.
IO Restriction.
To mitigate harmful activities such as unauthorized data access or data exfiltration, the sandbox imposes strict input/output (IO) restrictions. These restrictions can include limitations on the size and type of files that can be read from or written to the disk and restrictions on the use of network sockets for sending or receiving data. By controlling the IO operations, the sandbox can prevent many common exploits and ensure that the code running within it cannot interfere with other processes or the host system.
Isolated File System.
The sandbox employs an isolated file system to provide a safe execution environment for the code. This means that the process running in the sandbox has its virtual file system, which is separated from the host’s file system. The isolated nature of this file system ensures that even if a process within the sandbox attempts to modify or delete files, these changes will not affect the host system or other sandboxes. It acts as a layer of security, protecting the host from potential threats and maintaining the integrity of the overall system.
System Libraries Redirection.
To maintain a consistent and controlled environment, the sandbox redirects calls to system libraries to sandbox-specific versions. This is done to prevent code from using certain functions directly from the host’s system libraries, which could result in unpredictable behavior or security vulnerabilities. The redirected libraries are often limited to a subset of functionalities deemed safe and necessary for executing programs within the sandbox, thus enforcing the security policies and ensuring that the running programs behave as expected.
Single-threaded Evaluation.
In the context of the sandbox environment, single-threaded evaluation is the process of executing code using only one thread of execution. This approach simplifies resource management and timing measurements, avoiding the complexities associated with multi-threaded execution, such as synchronization, race conditions, and deadlocks. Single-threading can be particularly important in educational or testing settings where reproducibility and fairness are critical, ensuring that every student or user’s code is evaluated with the same computational resources.
Code Efficiency Measurement.
Figure 5 shows the overview of the code execution pipeline. We gauge the Solution Instantiation and Test Ease Evaluation time spans as the execution runtime.
Figure 5: Sandbox Execution Pipeline. 1) Test Case Generation. We first employ the corresponding test case generator for each task to produce a comprehensive set of test cases for the subsequent evaluation. 2) Context Initialization. To prevent any unexpected code behavior, the sandbox environment is meticulously reinitialized for each new task. This phase ensures that all the common libraries required for executing the solution are loaded. 3) Solution Instantiation. The solution under evaluation will be encapsulated as a solution class. 4) Test Case Evaluation. Each test case the generator provides will be rigorously executed against the solution. A solution must successfully pass all the test cases to be deemed valid. 5) Clean up. The final stage involves the sandbox dutifully clearing the namespace environment and the temporary directory. Mercury records the time consumed during the stage of Solution instantiation and Test Ease Evaluation as the primary metric for assessing computational efficiency.
A.4 DPO Experiment Details
Dataset Construction.
For every task problem $T^{i}$ in Mercury, we randomly selected two solutions from the task solution set $\{s^{i}_{w},s^{i}_{l}\}\sim T^{i}_{solution}$, to construct the preference dataset $D=\{P^{i},s^{i}_{w},s^{i}_{l}\}$, where $p^{i}$ is the prompt, $s^{i}_{w}$ has a faster runtime than $s^{i}_{l}$.
Model Initialization.
RLHF [45] typically begins with a reference LLM $\pi_{ref}$. Here, we initialize $\pi_{ref}$ by maximizing the likelihood of faster code completions $(p,s_{w})\sim D$, so that $\pi_{ref}=\arg\max_{\pi}E_{(p,s_{w})\sim D}\left[\log\pi(s_{w}|p)\right]$. This procedure helps mitigate the distribution shift between the true reference distribution and $\pi_{ref}$.
Optimization.
We optimize the target LLM $\pi_{\theta}$ to minimize $\mathcal{L}_{DPO}$ for the given $\pi_{ref}$ and $D$ and desired hyperparameter $\beta$. The gradient with respect to the parameters $\theta$ can be written as $\nabla_{\theta}\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})$.
$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})=-E_{(x,s_{w},s_{l})\sim D}\left[\log\alpha(\beta\log\frac{\pi_{\theta}(s_{w}|p)}{\pi_{ref}(s_{w}|p)})-\log\frac{\pi_{\theta}(s_{l}|p)}{\pi_{ref}(s_{l}|p)})\right]$ (2)
$\nabla_{\theta}\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})=\\ -\beta E_{(p,s_{w},s_{l})\sim D}\left[\underbrace{\alpha(\hat{r}_{\theta}(p,s_{l})-\hat{r}_{\theta}(p,s_{w}))}_{\textit{higher weight for wrong estimate}}\left[\underbrace{\nabla_{\theta}\log\pi(s_{w}|p)}_{\textit{increase likelihood of $s_{w}$}}-\underbrace{\nabla_{\theta}\log\pi(s_{l}|p)}_{\textit{decrease likelihood of $s_{l}$}}\right]\right]$ (3)
Intuitively, the gradient of the loss function $\mathcal{L}_{DPO}$ increases the likelihood of the preferred completions $s_{w}$ and decreases the likelihood of dis-preferred completions $s_{l}$, which are weighed by how much higher the implicit reward model $\hat{r}_{\theta}$ rates the dis-preferred completions, scaled by $\beta$, i.e., how incorrectly the implicit reward model orders the completions, accounting for the strength of the KL constraint.
A.5 External Libraries Utilized in Mercury
Raw LeetCode solutions typically commence without importing shared libraries. To avoid solution failure due to absent libraries, we proactively import the libraries listed in Figure 6 during the sandbox Context Initialization phase. Note that all these libraries are imported in a temporary namespace of which the sandbox controls code behaviors.
Figure 6: External Libraries Imported in Mercury Evaluate Framework.
A.6 Model Details
Table 7: Model Details. We evaluated LLMs ranging from 1.3B to 34B.
Model Name Model Scale Link
deepseek-coder-1.3b-base 1.3B https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-base
starcoder2-3b 3B https://huggingface.co/bigcode/starcoder2-3b
deepseek-coder-6.7b-base 6.7B https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base
starcoder2-7b 7B https://huggingface.co/bigcode/starcoder2-7b
CodeLlama-7b-hf 7B https://huggingface.co/codellama/CodeLlama-7b-hf
CodeQwen1.5-7B 7B https://huggingface.co/Qwen/CodeQwen1.5-7B
CodeLlama-13b-hf 13B https://huggingface.co/codellama/CodeLlama-13b-hf
starcoder2-15b 15B https://huggingface.co/bigcode/starcoder2-15b
deepseek-coder-33b-base 33B https://huggingface.co/deepseek-ai/deepseek-coder-33b-base
CodeLlama-34b-hf 34B https://huggingface.co/codellama/CodeLlama-34b-hf
A.7 A Mercury Example
Figure 7: This case is drawn from the Mercury-eval benchmark. The upper block presents the problem statement with its example, while the subsequent portion exhibits the corresponding solutions. Although all solutions provided are functionally correct, they exhibit significant differences in runtimes.
A.8 Hardware-agnostic Evaluation
Figure 8: $Beyond$ scores of ‘deepseek-coder-33b’ (solid line) and ‘deepseek-coder-6.7b’ (dashed line) across varied Intel Skylake CPU configurations. The results show that $Beyond$ can keep consistent across different hardware configurations.
A.9 Prompts for Code Generation
To guarantee a fair comparison, we apply a unified one-shot prompt template for each pre-trained Code-LLM. As displayed in Figure 9, the prompt template contains one shot example as well as three placeholders: <task_content>, <code_starter>, and <code_completion>.
Figure 9: Code Generation Prompts. Lines 1 to 40 are the one-shot example. In Mercury experiments, we feed the pretty_content field to the placeholder <task_content>, the prompt field to the placeholder <code_starter>, and the solution field to the placeholder <code_completion>
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and precede the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [TODO]
[…]
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [TODO]
[…]
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [TODO]
[…]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [TODO]
[…]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [TODO]
[…]
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [TODO]
[…]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [TODO]
[…]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [TODO]
[…]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [TODO]
[…]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [TODO]
[…]
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [TODO]
[…]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [TODO]
[…]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [TODO]
[…]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [TODO]
[…]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [TODO]
[…]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
-Mercury: A Computational Efficiency Benchmark for Code Large Language Models
+Mercury: A Code Efficiency Benchmark for Code Large Language Models
 Abstract
-Amidst the recent strides in evaluating Large Language Models for Code (Code-LLMs), existing benchmarks have mainly focused on functional correctness, overlooking the importance of computational efficiency.
-To fill the gap, we present Mercury, the first computational efficiency benchmark for Code-LLMs.
-It comprises 1,889 Python tasks, each with adequate solutions to support a runtime distribution.
-Based on the distribution, we introduce a new metric $Beyond$, which computes a runtime-percentile-weighted $Pass$ score to reflect functional correctness and computational efficiency simultaneously.
-On Mercury, leading Code-LLMs can achieve 67% on $Pass$, while less then 50% on $Beyond$.
-Given that an ideal $Beyond$ score would be aligned with the $Pass$ score, it indicates that while Code-LLMs exhibit impressive capabilities in generating functionally correct code, there remains a notable gap in their efficiency.
-Finally, our empirical experiments reveal that Direct Preference Optimization (DPO) serves as a robust baseline for enhancing computational efficiency compared with Supervised Fine Tuning (SFT), which paves a promising avenue for future exploration of efficient code generation.
-[^1] Our code and data are available on GitHub: https://github.com/Elfsong/Mercury..
+Amidst the recent strides in evaluating Large Language Models for Code (Code LLMs), existing benchmarks have mainly focused on the functional correctness of generated code, neglecting the importance of their computational efficiency.
+To fill the gap, we present Mercury, the first code efficiency benchmark for Code LLMs.
+It comprises 1,889 Python tasks, each accompanied by adequate solutions that serve as real-world efficiency baselines, enabling a comprehensive analysis of the runtime distribution.
+Based on the distribution, we introduce a new metric Beyond, which computes a runtime-percentile-weighted Pass score to reflect functional correctness and code efficiency simultaneously.
+On Mercury, leading Code LLMs can achieve 65% on Pass, while less than 50% on Beyond.
+Given that an ideal Beyond score would be aligned with the Pass score, it indicates that while Code LLMs exhibit impressive capabilities in generating functionally correct code, there remains a notable gap in their efficiency.
+Finally, our empirical experiments reveal that Direct Preference Optimization (DPO) serves as a robust baseline for enhancing code efficiency compared with Supervised Fine Tuning (SFT), which paves a promising avenue for future exploration of efficient code generation.
+[^1] Our code and data are available on GitHub: https://github.com/Elfsong/Mercury.
 1 Introduction
-The domain of code generation, which aims to empower computers to autonomously generate code based on natural language task instructions (NL2Code), constitutes a longstanding and fascinating research objective [42].
-The recent emergence of Large Language Models (LLMs) has spurred the development of a new wave of code generation models [40, 31, 35, 11, 27, 4], which leverage the impressive language understanding and generative capabilities of LLMs to drive forward the ambitious goal of producing high-quality and machine-generated code.
-In parallel with the model development, recent code generation benchmarks mainly focus on evaluating functional correctness in the fuzz testing style [25].
-This testing approach assesses the outcome congruence between the generated and anticipated solutions by executing bespoken test cases.
-For instance, HumanEval [8] and MBPP [3] collected a small but high-quality set of handcrafted tasks, and EvalPlus [24] further consolidates these two above benchmarks by augmenting the test scope.
+The domain of code generation, which aims to empower computers to autonomously generate code based on natural language task descriptions (NL2Code), has long been considered a promising way to facilitate interaction between humans and computers [46, 44].
+The recent emergence of Large Language Models (LLMs) has spurred a new wave of NL2Code models [34, 38, 12, 29, 4], which leverage the impressive language understanding and generative capabilities of LLMs to drive forward the ambitious goal of synthesizing high-quality code from natural language instructions.
+To measure the quality of code, recent code generation benchmarks mainly focus on evaluating their functional correctness via test case fuzzing [27].
+This approach assesses the outcome congruence between the LLM-generated and canonical solutions by executing bespoken test cases.
+For instance, HumanEval [8] and MBPP [3] collected a small but fine set of handcrafted tasks with test cases.
+EvalPlus [26] further consolidates these two above benchmarks by augmenting the case scope.
 On the contrary, APPS [15] widely gathered over 5,000 public coding tasks from online platforms.
-Despite these strides, there is a discernible oversight in current code generation benchmarks concerning the computational efficiency evaluation, although that is critical in real-world software development [43, 44].
-Moreover, for most tasks in the above benchmarks, handcrafting diverse solutions and comprehensive test cases to cover all possible scenarios is a time-consuming endeavor [24].
-In light of these findings, we highlight several vital limitations inherent in the existing code generation benchmarks:
-Figure 1: While both LLM-generated codes successfully follow the task instruction, the right snippet notably excels in computational efficiency, executing in a mere 121 ms compared to the 5,714 ms consumed by the left snippet.
-On the premise of ensuring functional correctness, optimizing computing efficiency is a ubiquitous objective in software development.
+Despite these strides, there is a discernible oversight in current code generation benchmarks concerning the code efficiency evaluation, although that is critical in software development [45, 47].
+Moreover, handcrafting diverse solutions and test cases to cover all scenarios is infeasible [26].
+In light of these findings, we highlight vital limitations inherent in the existing code generation benchmarks:
+Figure 1: Executing these two LLM-generated codes on 100 test cases.
+While both codes successfully follow the task instruction and pass all test cases, the right snippet notably excels in code efficiency, completing in a mere 121 ms compared to the 5,714 ms consumed by the left snippet.
+As Code LLMs become widely used in the real world, code efficiency determines factual productivity, where Mercury can gauge the vital metric.
   1. 1.
-    Absence of Computational Efficiency Evaluation.
-Existing code generation benchmarks focus on assessing functional correctness while overlooking the evaluation of computational efficiency [8, 3, 14].
+    Absence of code Efficiency Evaluation.
+Existing code generation benchmarks focus on assessing functional correctness while overlooking the evaluation of code efficiency [8, 3, 15].
 As illustrated in Figure 1, despite both code snippets can handle the sorting task functionally, the right efficient solution (121 ms) is nearly 50 times faster than the left inefficient solution (5,714 ms).
-This striking runtime differentiation underscores the necessity of incorporating computational efficiency assessments within code generation benchmarks to ensure that synthesized code is not only correct but also efficient.
+This striking runtime differentiation underscores the necessity of incorporating code efficiency assessments within code generation benchmarks, encouraging Code LLMs to produce not only correct but also efficient code.
   2. 2.
     Insufficient Test Case Coverage.
-Most code generation benchmarks obtain a limited set of handcrafted test cases for assessing functional correctness, potentially overlooking edge cases and nuanced code behaviors [8, 3].
+As shown in Table 1, most code generation benchmarks manually build a small number of test cases or extract the accompanying test cases from existing resources, potentially overlooking edge cases and nuanced code behaviors [8, 3].
 For example, Figure 8 displays that HumanEval #55 contains only 3 test cases, testing up to the 12th Fibonacci number [8].
 Its given canonical solution will quickly reach the recursion depth limitation when computing a larger Fibonacci number (the recursion limitation depends on the environment).
-Therefore, notwithstanding the generated code satisfies all test cases, such success does not necessarily equate to assurance of functional correctness and much less to computational efficiency.
+Therefore, notwithstanding the generated code satisfies all test cases, such success does not necessarily equate to assurance of functional correctness and much less to code efficiency.
   3. 3.
     Lack of Task Diversity.
-Another noticeable deficit of existing code generation benchmarks is the insufficient diversity and complexity in their constituent tasks [8, 3, 25].
-Since most benchmarks only consist of elementary-level problems, recent Code-LLMs can effortlessly tackle most tasks regardless of their actual capacities [44].
-This flaw results in these benchmarks failing to pose a substantial challenge to Code-LLMs and truly reflect their underlying potential.
-Computational Efficiency.
-Computational efficiency refers to the performance measure of time and space complexity to accomplish a specific task.
+Another noticeable deficit of existing code generation benchmarks is the insufficient diversity and complexity in their tasks [8, 3, 27].
+Since most benchmarks only consist of elementary-level programming tasks, recent Code LLMs can effortlessly tackle most tasks regardless of their actual capacities [47].
+This flaw results in these benchmarks failing to pose a substantial challenge to Code LLMs and truly reflect their underlying potential.
+Code Efficiency.
+Code efficiency refers to the performance measure of time and space complexity to accomplish a specific task.
 Efficient code can improve user experience, save energy, and make applications more sustainable and cost-effective.
-Compared with the scalable memory space, execution time is the performance bottleneck of most programs.
-Consequently, this work focuses on the time dimension of computational efficiency.
+Compared with the scalable memory space, execution time is the performance bottleneck of most codes.
+Consequently, this work focuses on the time dimension of code efficiency.
 Figure 2: An overview of Mercury dataset.
 Each Mercury task has a task description, a test case generator, a prompt & entry point, and corresponding solutions.
-To evaluate computational efficiency, we introduce the $Beyond$ metric, which signifies the runtime percentile of the LLM-generated code on the runtime distribution supported by corresponding solutions.
+To evaluate code efficiency, we introduce the Beyond metric, which signifies the runtime percentile of the LLM-generated code on the runtime distribution supported by corresponding solutions.
 In this example, the LLM-generated code executes in 521 ms, outpacing 86.18% of collected solutions on the runtime distribution.
@@
 Our Benchmark.
-In this work, we introduce Mercury, a novel code generation benchmark designed to assess computational efficiency.
-As depicted in Figure 2, Mercury comprises 1,889 Python programming tasks with three difficulty stratification, which is divided into two datasets for model fine-tuning and evaluation separately.
-For each task, we assign a test case generator to remedy the shortfall of test case coverage.
-In measuring computational efficiency, the primary challenge stems from normalizing the absolute runtime across tasks that have diverse runtime ranges.
-Thus we collect and locally execute numerous solutions for each task to form a runtime distribution and leverage the runtime percentile of LLM-generated code on the distribution instead of the absolute runtime to evaluate computational efficiency.
+In this work, we introduce Mercury, a novel code generation benchmark designed to assess and improve the code efficiency of Code LLMs.
+As depicted in Figure 2, Mercury comprises 1,889 Python programming tasks with three difficulty stratification, which is divided into two datasets for model evaluation and fine-tuning separately.
+For each evaluation task, we assign a test case generator to remedy the shortfall of test case coverage.
+In measuring code efficiency, the primary challenge stems from normalizing the absolute runtime across tasks that have diverse runtime ranges.
+Thus, we collect and locally execute numerous historical solutions for each task to form a runtime distribution and leverage the runtime percentile of LLM-generated code on the distribution instead of the absolute runtime to evaluate code efficiency.
 Furthermore, to mitigate performance discrepancies attributed to irrelevant processes and diverse hardware configurations, we set up an isolated sandbox environment for task execution to establish local runtime distributions.
-Contribution.
-Our work aimed to fill the computational efficiency evaluation gap in code generation benchmarks with the following key contributions:
+Contribution. Our work aimed to fill the code efficiency evaluation gap in code generation benchmarks with the following key contributions:
   • Dataset.
-We introduce a novel code generation dataset Mercury designed to assess and improve computational efficiency in Section 2, accompanied by an extensible open-source data collection framework for enriching Mercury with more tasks and languages.
-  • Metric. We propose the first efficiency-focused code generation metric $Beyond$ and establish a benchmark to evaluate leading Code-LLMs using this metric in Section 3.
+We collect a novel code generation dataset Mercury designed to assess and improve Code LLM code efficiency in Section 2, accompanied by an extensible open-source data collection framework for enriching Mercury with more tasks and programming languages.
+  • Metric. We propose the first efficiency-focused code generation metric Beyond and establish a benchmark to evaluate leading Code LLMs using this metric in Section 3.
   • Baselines.
-In Section 4, we detail our extensive analysis of baselines to enhance computational efficiency while maintaining functional correctness.
-Experiment results reveal that despite Code-LLMs excelling in functional correctness, there is still considerable potential to elevate efficiency.
-Table 1: A comparison of Mercury to existing NL2Code benchmarks. Mercury distinguishes itself by including a set of distilled high-quality solutions and a dedicated test case generator for each task.
+In Section 4, we detail our extensive analysis of two baselines to enhance code efficiency while maintaining functional correctness.
+Experiment results reveal that despite Code LLMs excelling in functional correctness, there is still considerable potential to elevate efficiency.
+Table 1: A comparison of Mercury to existing NL2Code benchmarks.
+Mercury distinguishes itself by including a set of distilled high-quality solutions and a dedicated test case generator for each task.
+* signifies that the solution number can be further expanded by the data collection framework.
 Benchmarks Tasks Sources Cases Solutions Difficulty
@@
 APPS 5,000 Online 21.2 23.4 3
-Mercury 256 Online + Filters $\infty$ 18.4 3
-Figure 3: An example of insufficient test cases. Even though the code passed all test cases in the dashed-line box, it remains vulnerable to timeout or stack overflow when subjected to a larger input.
+Mercury 256 Online + Filters $\infty$ 18.4 * 3
 2 Mercury Datasets
-We initiate this work by collecting public programming tasks on Leetcode [^2] https://leetcode.com/problemset/algorithms/.
+We initiate this work by collecting public programming tasks on Leetcode [23].
 Subjecting these questions to a series of filters, we distilled them down to 1,889 high-quality tasks.
@@
 We delineate the definition of Mercury fields in the Appendix Table 5.
-Task Filters.
-Mercury tasks originate from public programming problems on Leetcode. To assure the quality and uniformity of the dataset, we distilled gathered tasks based on the following conditions:
+Task Filters. Mercury tasks originate from public programming problems on Leetcode. To assure the quality and uniformity of the dataset, we distilled gathered tasks based on the following conditions:
+  1. 1.
     Number of Solutions.
@@
 After excluding these tasks, Mercury tasks possess an average of 18.4 unique solutions.
+  2. 2.
     Restricted Data Structure.
-Above the inherent Python data types, Mercury also incorporates two custom data types: Binary Tree and Linked List (the specific structure definitions can be found in Figure 4), which increases Mercury’s diversity and escalates its difficulty level.
+Above the inherent Python data types, Mercury also incorporates two custom data types: Binary Tree and Linked List (the specific structure definitions can be found in Appendix Figure 4), which increases Mercury’s diversity and escalates its difficulty level.
 Tasks that contain other data structures will be removed.
+  3. 3.
     Unique Outputs.
@@
 Evaluating all possible answers can drastically complicate the test case verification process.
-To eliminate this problem, we harness the corresponding test case generator to generate $N$ test cases $T_{i}=\langle Input_{i},Output_{i}\rangle\ s.t.\ i\in\{0,1,\cdots,N\}$ and execute $T$ on different solutions $S_{m}\ s.t.\ m\in\{0,1,\cdots,M\}$ to observe if all outputs $Output_{i}=S_{m}(Input_{i})$ remain identical.
+To eliminate this problem, we harness the corresponding test case generator to generate $N$ test cases $T_{i}=\langle Input_{i},Output_{i}\rangle\ s.t.\ i\in\{0,1,\cdots,N\}$ and execute $T$ on different solutions $S_{m}\ s.t.\ m\in\{0,1,\cdots,M\}$ to observe if all $Output_{i}=S_{m}(Input_{i})\ s.t.\ i\in\{0,1,\cdots,N\}$ remain identical.
 Any tasks that potentially yield non-unique answers were subsequently excluded.
@@
 To alleviate this issue, Mercury inherits the difficulty categorization from Leetcode, i.e., Easy, Medium, and Hard.
-The stratification aims to probe the upper bounds of Code-LLM capabilities, delivering a more evident distinction between various Code-LLMs.
+The stratification aims to probe the upper bounds of Code LLM capabilities, delivering a more evident distinction between various Code LLMs.
 Test Case Generator.
 Manual creation of test cases can be a laborious process.
-To gather sufficient test cases to conduct an exhaustive assessment, we assign a test case generator for each task, which can produce a full range of test cases to thoroughly evaluate the functional correctness and efficiency of given solutions.
-Specifically, We reformulate the task description as plain text (see $pretty\_content$ in Table 5) and then feed it into GPT-4 [31] to generate an initial test case generator snippet.
-To confirm the effectiveness of the initial generator, we subsequently create a batch of test cases (24 cases in our setting) by the generator and submit them to the Leetcode online judge system.
-Should any of the generated test cases not pass the Leetcode validation, we manually revise the generator until all generated cases can be successfully validated.
-3 Computational Efficiency Metric
-In the domain of software development, computational efficiency can be defined as the absolute code runtime for executing a given test case set [7].
-Nonetheless, a primary obstacle in benchmarking computational efficiency is normalizing runtime measurements across disparate environments.
+To gather sufficient test cases to conduct an exhaustive assessment, we assign a test case generator for each evaluation task, which can produce a full range of test cases to thoroughly evaluate the functional correctness and code efficiency of given solutions.
+Specifically, We feed $pretty\_content$ into GPT-4 [34] to generate an initial test case generator snippet.
+To confirm the effectiveness of the initial generator, we subsequently create 24 test cases by the generator and submit these cases to the Leetcode Online Judge (OJ) system.
+Should any of the generated test cases not pass the LeetCode OJ validation, we manually revise the generator until all generated cases can be successfully validated.
+3 Code Efficiency Metric
+In the domain of software development, code efficiency can be defined as the absolute code runtime for executing a given test case set [7].
+Nonetheless, a primary obstacle in benchmarking code efficiency is normalizing runtime measurements across disparate environments.
 For instance, a sub-optimal solution might have a faster absolute runtime on high-performance hardware than an optimal solution on low-performance hardware.
 Moreover, different operation systems and code interpreters may also fluctuate the code runtime.
-Therefore, absolute runtime fails as a consistent and reliable computational efficiency benchmark metric.
+Therefore, absolute runtime fails as a consistent and reliable code efficiency benchmark metric.
 To address this issue, an intuitive approach involves modeling a devoted runtime distribution for each task and calculating the average runtime percentiles of LLM solution samples over the runtime distribution.
-With this idea in mind, we proposed a normalized computational efficiency metric $Beyond$:
+With this idea in mind, we proposed a normalized code efficiency metric Beyond:
 $p_{k}^{n}=\frac{max(R^{n})-clip(r_{k}^{n},min(R^{n}),max(R^{n}))}{max(R^{n})-min(R^{n})},\qquad Beyond=\frac{\sum\nolimits_{N,K}^{n=0,k=0}p_{k}^{n}}{N\cdot K}.$ (1)
 Where $N$ is the total number of tasks, and $K$ denotes the size of LLM solution samples.
-For a specific task $n\in N$, $R^{n}$ is the runtime array corresponding to the collected solutions, and $r_{k}^{n}$ denotes the runtime for the $k$-th LLM solution.
-Runtime is defined as the period from the solution instantiation to the evaluation across all test cases, culminating with a successful termination (More engineering details can be found in Appendix Figure 5).
-Since any case failure of the $k$-th solution results in $r_{k}^{n}\rightarrow+\infty$ and then $p_{k}^{n}=0$, $Beyond$ reflects functional correctness as well.
+For a specific task $n\in N$, $R^{n}$ is the runtime array corresponding to the collected historical solutions, and $r_{k}^{n}\ s.t.\ k\in K$ denotes the runtime for the $k$-th LLM solution.
+$clip$ is a function to constraint the value $r_{k}^{n}$ in the range $[min(R^{n}),max(R^{n})]$.
+Runtime is defined as the period from the solution instantiation to the evaluation across all test cases, culminating with a successful termination (More engineering details can be found in Appendix Section A.3).
+Since any case failure of the $k$-th solution results in $r_{k}^{n}\rightarrow+\infty$ and then $p_{k}^{n}=0$, Beyond can reflect functional correctness as well.
 Untrusted Code Execution.
-Since most Code-LLMs are trained on an extensive code corpus from unverified sources, there is an intrinsic risk that these models may produce malicious code when driven by specific meticulous prompts [8].
+Since most Code LLMs are trained on an extensive code corpus from unverified sources, there is an intrinsic risk that these models may produce malicious code when driven by specific meticulous prompts [8].
 The direct execution of synthesized code raises significant security concerns.
@@
 Sandbox details are deliberated in Appendix A.3
-Hardware-agnostic Evaluation.
-To ensure fair comparison across diverse hardware configurations, we run each task $n$ with corresponding test cases locally and aggregate their runtimes into the runtime array $R^{n}$.
+Environment-agnostic Evaluation.
+To ensure fair comparison across diverse configurations, we run each task $n$ with corresponding test cases locally and aggregate their runtimes into the runtime array $R^{n}$.
 Appendix Figure 10 illustrates the Beyond score of two LLMs (‘deepseek-coder-33b’ and ‘deepseek-coder-6.7b’) over three distinct hardware specifications: the micro-tier (0.25 CPU cores), the small-tier (0.5 CPU cores), and the standard-tier (1 CPU core).
@@
 4 Experiments
-In this section, we present a series of baseline experiments to improve computational efficiency by training on Mercury-train dataset and assessing on the Mercury-eval dataset.
-Our empirical study encompasses 10 open-source LLMs a broad parameter spectrum from 1.3 to 34 billion.
-For each LLM, we compare the performance of the original model and two optimization strategies: Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), for their potential to optimize LLM generating functionally correct and computationally efficient code.
+In this section, we present a series of baseline experiments to improve code efficiency by training on Mercury-train dataset and assessing on the Mercury-eval dataset.
+Our empirical study encompasses 10 open-source LLMs with a broad parameter spectrum from 1.3 to 34 billion.
+For each LLM, we compare the performance of the original model and two optimization strategies, Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), for their potential to optimize LLM generating functionally correct and computationally efficient code.
 Finally, we analyzed the underlying factors contributing to the failure of LLMs on the Mercury-eval dataset.
@@
 Within the SFT [6] method, an LLM undergoes additional training on a small dataset, which aims to specialize the LLM to perform better on certain tasks correlated to the training dataset.
-To optimize the computational efficiency performance of Code-LLMs, the most intuitive strategy is to fine-tune the Code-LLM using optimal runtime solutions.
-In our experimental setup, we apply the unified prompt template for each Code-LLM to ensure a fair comparison.
+To optimize the code efficiency performance of Code LLMs, the most intuitive strategy is to fine-tune the Code LLM using optimal runtime solutions.
+In our experimental setup, we apply a unified prompt template for each Code LLM to ensure a fair comparison.
 The “pretty_content” attribute fills the <task_content> placeholder, the “prompt” attribute fills the <code_starter> placeholder, and the <code_completion> placeholder is completed with the fastest solutions.
-To steer Code-LLMs towards generating the intended code completion, we prepend a one-shot example to the prompt template.
+To steer Code LLMs towards generating the intended code completion format, we prepend a one-shot example to the prompt template.
 Appendix Figure 9 presents the prompt template.
@@
 To enable LLMs to align with human preferences while preserving their functional capabilities, existing methodologies employ reinforcement learning with human preference feedback (RLHF).
-However, RLHF introduces additional complexities and potential instabilities, necessitating significant computing resources and additional reward model training [45, 5, 37].
+However, RLHF introduces additional model complexities and potential instabilities, necessitating significant computing resources and extra reward model training [48, 5, 40].
 DPO [36] bypasses these challenges by explicitly mapping reward functions and the optimal objective.
 This connection demonstrates that maximizing rewards under specific constraints can be effectively addressed through a singular training phase based on data reflecting human preferences.
-Consequently, DPO provides a stable, efficient, and computation-friendly alternative, removing the necessity for separate reward modeling.
-The DPO dataset construction and training procedure are elaborated in Appendix Section A.4.
+The DPO training procedure is elaborated in Appendix Section A.4.
 4.2 Functional Correctness Benchmarks
@@
 Each problem in this dataset consists of three components: a task description, an associated code solution, and three automated test cases to validate the code functionality [3].
-Both HumanEval and MBPP harness the metric $Pass$ to measure the Code-LLMs’ functional correctness, where a task is considered solved if it passes all test cases, and the total fraction of solved passed is reported [20].
+Both HumanEval and MBPP harness the metric Pass to measure the Code LLMs’ functional correctness, where a task is considered solved if the given solution passes all test cases, and the total fraction of solved tasks is reported as $Pass=N_{solved}/N_{total}$ [21].
 4.3 Experimental Setups
-Configuration.
 […]
@@
 4.4 Empirical Results
-Functional correctness is the prerequisite for evaluating computational efficiency for the code generation task.
-Our primary objective is to enhance the computational efficiency without compromising the functional correctness.
-To this end, we first introduce the existing metric $Pass$ to gauge the functional correctness [8] and then leverage $Beyond$ to provide a holistic evaluation, encompassing both computational efficiency and functional correctness.
-Finally, we measure the $Gap$ between $Beyond$ and $Pass$ to indicate the baseline effectiveness.
-These experiments aim to scrutinize the innate capabilities of cutting-edge Code-LLMs and their potential after baseline fine-tuning.
+Functional correctness is the prerequisite for evaluating code efficiency for the code generation task.
+Our primary objective is to enhance the code efficiency without compromising the functional correctness.
+To this end, we first introduce the existing metric Pass to gauge the functional correctness [8] and then leverage Beyond to provide a holistic evaluation, encompassing both code efficiency and functional correctness.
+Finally, we measure the Gap between Beyond and Pass to mirror the baseline ability for improving efficiency while preserving correctness.
+These experiments aim to investigate the innate capabilities of cutting-edge Code LLMs and their potential after baseline fine-tuning.
 Therefore, extensive parameter optimization and prompt engineering were not pursued.
 To deliver a comprehensive evaluation, we have further integrated the HumanEval and MBPP benchmarks as supplementary measures for appraising functional correctness [8, 3].
+Figure 3: The horizontal axis represents the score for functional correctness, while the vertical axis indicates the score for code efficiency.
+The left figure illustrates the performance of the baseline model, whereas the right one depicts the performance after DPO tuning.
+Model points located nearer to the diagonal line exhibit a more equitable balance between functionality and efficiency.
 Functional Correctness.
-Table 2 lists functional correctness benchmarks over various models, showing that larger models tend to provide better functional correctness.
-Except for the smallest model “deepseek-coder-1.3b-base”, DPO consistently enhances the overall $Pass$ scores across most Code-LLMs, while SFT diminishes functional correctness on the largest two Code-LLMs.
-These findings suggest that smaller models may struggle to integrate new knowledge while preserving their original functionality, and SFT may induce catastrophic forgetting in larger models in the pursuit of heightened computational efficiency.
-Moreover, it is observable that the functional correctness metrics of each model consistently diminish as the difficulty level increases, where Mercury difficulty stratification probes the model’s upper capability better than the other two benchmarks.
-Table 2: Functional correctness ($Pass$) evaluation results over LLMs.
-The underlined values denote the top-performed approaches among the original model and the established baselines.
+Table 2 lists Pass scores over various Code LLMs, showing that larger models tend to provide better functional correctness.
+Except for the smallest model “deepseek-coder-1.3b-base”, DPO invariably enhances the overall Pass scores across most Code LLMs, while SFT diminishes functional correctness on the largest two Code LLMs.
+These findings suggest that smaller models may struggle to integrate new knowledge while preserving their original functionality, and SFT may induce catastrophic forgetting in the pursuit of heightened code efficiency.
+Moreover, it is evident on Mercury that Pass scores of each model consistently decline as the difficulty level increases, indicating that the Mercury difficulty stratification is effective at probing the upper limitation of each Code LLM compared to the auxiliary benchmarks.
+Table 2: Functional correctness (Pass) evaluation results.
+The underlined values denote the top-performed approaches among the original model and baselines.
 The bolded values denote the best performance on each benchmark.
+We sample one solution for each task to calculate pass score.
 Mercury
 Model Name HumanEval MBPP Easy Medium Hard Overall
-[…]
-[…]
-[…]
+deepseek-coder-1.3b-base 28.7 55.4 60.7 52.8 23.2 38.1
++ SFT 24.2 46.2 58.9 53.6 25.3 36.2 (-1.9)
++ DPO 29.1 50.2 61.4 53.6 20.0 35.9 (-2.2)
 starcoder2-3b 31.7 57.4 56.1 52.1 21.6 37.8
-[…]
-[…]
++ SFT 29.0 47.2 60.7 58.8 25.3 38.8 (+1.0)
++ DPO 33.5 59.6 62.5 61.0 23.4 41.1 (+3.3)
 deepseek-coder-6.7b-base 47.6 70.2 69.3 68.9 56.1 61.0
-[…]
-[…]
++ SFT 56.1 59.6 69.1 71.4 57.7 62.2 (+1.2)
++ DPO 54.3 72.8 74.1 72.6 58.9 65.4 (+4.4)
 starcoder2-7b 35.2 54.4 63.6 61.7 29.2 44.3
 + SFT 42.9 57.2 64.8 58.5 31.3 47.5 (+3.2)
-[…]
-[…]
++ DPO 55.4 61.4 74.8 66.9 32.6 53.6 (+9.3)
+CodeLlama-7b-hf 33.5 52.0 55.7 41.7 12.9 29.6
 + SFT 29.5 47.6 58.9 38.5 16.1 31.3 (+1.7)
-[…]
-[…]
-[…]
-[…]
++ DPO 38.7 49.2 67.5 45.7 17.9 36.1 (+6.5)
+CodeQwen1.5-7B 51.8 72.2 70.0 70.1 49.7 61.1
++ SFT 54.3 74.8 70.9 67.9 49.7 61.9 (+0.8)
++ DPO 55.5 75.4 72.5 66.9 45.7 61.1 (+0)
 starcoder2-15b 46.3 66.2 69.5 65.4 50.3 58.0
 + SFT 51.6 69.2 72.0 68.9 51.7 61.3 (+3.3)
-[…]
-[…]
++ DPO 57.0 72.8 78.0 73.8 54.7 66.7 (+8.7
+CodeLlama-13b-hf 37.8 62.4 76.8 60.5 18.4 39.6
 + SFT 39.5 59.8 65.5 54.8 19.5 39.5 (-0.1)
-[…]
++ DPO 49.1 64.4 78.6 60.0 29.0 50.1 (+10.6)
 deepseek-coder-33b-base 54.3 73.2 70.9 67.9 62.3 65.0
 + SFT 58.1 74.8 61.8 58.0 47.1 58.7 (-6.3)
-[…]
++ DPO 72.9 80.6 78.9 76.5 61.6 73.4 (+8.4)
 CodeLlama-34b-hf 48.2 65.4 77.7 63.7 32.4 52.4
 + SFT 52.8 68.2 61.8 58.0 26.2 47.5 (-4.9)
-[…]
-Code Efficiency.
-Table 3: Computational efficiency ($Beyond$) evaluation results across three difficulty levels.
++ DPO 65.9 75.2 83.9 68.4 63.2 70.6 (+18.2)
+Table 3: Code efficiency (Beyond) evaluation results across three difficulty levels.
 The bolded value indicates the top performance for each metric, while the underlined values denote the most effective approaches among the original model and the baselines.
+In our experiment, we sample 5 solutions for each task to calculate Beyond score.
 Model name Easy Medium Hard Overall Gap
-[…]
+deepseek-coder-1.3b-base 47.97 39.77 19.26 35.62 9.85
 + SFT 42.58 38.12 18.67 33.04 (-2.58) 12.74 (+2.89)
-[…]
-[…]
-[…]
-+ DPO 43.70 41.02 12.99 32.42 (-0.99) 16.33 (6.61)
-[…]
-+ SFT 51.37 52.71 44.28 49.39 (+1.09) 16.55 (0.16)
-[…]
-[…]
-[…]
-[…]
-[…]
-+ SFT 39.75 26.89 9.55 25.41 (-2.04) 12.48 (3.21)
-[…]
-[…]
++ DPO 46.91 42.27 16.78 35.21 (-0.41) 9.64 (-0.21)
+starcoder2-3b 43.55 41.91 15.21 33.40 9.72
++ SFT 44.64 42.10 15.72 34.01 (+0.61) 14.04 (+4.31)
++ DPO 43.70 41.02 12.99 32.42 (-0.99) 16.33 (+6.61)
+deepseek-coder-6.7b-base 48.80 51.16 45.11 48.29 16.40
++ SFT 51.37 52.71 44.28 49.39 (+1.09) 16.55 (+0.16)
++ DPO 56.25 52.35 40.62 49.70 (+1.41) 18.73 (+2.34)
+starcoder2-7b 50.23 51.29 20.25 40.37 10.95
++ SFT 42.21 44.02 21.09 35.61 (-4.77) 15.80 (+4.84)
++ DPO 53.52 51.41 17.35 40.56 (+0.18) 17.41 (+6.46)
+CodeLlama-7b-hf 42.55 30.99 8.88 27.45 9.27
++ SFT 39.75 26.89 9.55 25.41 (-2.04) 12.48 (+3.21)
++ DPO 54.14 34.48 11.10 33.29 (+5.84) 10.46 (1.19)
+CodeQwen1.5-7B 51.11 53.56 39.03 47.78 15.35
 + SFT 54.16 51.43 38.05 47.82 (0.04) 14.91 (-0.43)
-[…]
++ DPO 56.07 51.55 38.05 48.52 (0.74) 13.12 (-2.22)
 starcoder2-15b 58.18 52.09 37.34 49.17 12.55
-+ SFT 53.54 52.77 37.73 47.92 (-1.25) 16.22 (3.67)
-[…]
++ SFT 53.54 52.77 37.73 47.92 (-1.25) 16.22 (+3.67)
++ DPO 68.29 59.54 48.97 58.95 (9.78) 10.81 (-1.74)
 CodeLlama-13b-hf 57.00 44.25 12.99 38.01 13.79
 + SFT 44.95 39.96 13.55 32.70 (-5.31) 13.78 (-0.01)
-[…]
++ DPO 67.09 55.72 19.72 47.39 (9.38) 8.47 (-5.32)
 deepseek-coder-33b-base 51.26 48.90 45.43 48.53 18.50
 + SFT 40.33 37.75 36.82 38.32 (-10.21) 17.30 (-1.20)
-[…]
++ DPO 74.59 68.91 55.98 66.47 (+17.94) 5.79 (-12.70)
 CodeLlama-34b-hf 56.28 48.21 22.96 42.40 15.49
 + SFT 45.49 44.96 20.73 36.91 (-5.50) 11.61 (-3.88)
-[…]
-Regarding the NL2Code task, once functional correctness has been assured, attention naturally pivots to enhancing computational efficiency.
-As depicted in Table 3, we investigate computational efficiency metrics across a spectrum of Code-LLMs.
-Experiments demonstrate that DPO yields a stable enhancement in computational efficiency from models exceeding 6.7 billion parameters.
-In contrast, SFT detracts most $Beyond$ scores from original models, suggesting that the plain SFT may not be a feasible strategy for enhancing computational efficiency.
++ DPO 78.55 60.95 51.94 63.94 (+21.54) 8.01 (-7.47)
+Code Efficiency.
+Regarding the NL2Code task, once functional correctness has been assured, attention naturally pivots to enhancing code efficiency.
+As depicted in Table 3, we investigate code efficiency metrics across a spectrum of Code LLMs.
+Experiments demonstrate that DPO yields a stable enhancement in code efficiency from models exceeding 6.7B parameters.
+In contrast, SFT detracts most Beyond scores from original models, suggesting that the plain SFT may not be a feasible strategy for enhancing code efficiency.
 Further analysis compares the Gap between Beyond and $Pass$.
-Since the ideal $Beyond$ should be aligned with $Pass$, it shows how much the baseline method shrinks the gap between functional correctness and computational efficiency.
-Our findings indicate that DPO substantially narrows $Gap$ in models larger than 15 billion parameters.
+Since the ideal Beyond should be aligned with Pass (where the LLM-generated solution is correct and faster than all historical solutions), it shows how much the baseline method shrinks the gap between functional correctness and code efficiency.
+Our findings indicate that DPO substantially narrows Gap in models larger than 15B parameters.
 However, Gap tends to widen in smaller models under the same configuration.
-This implies that larger models possess a greater capacity to assimilate the nuanced knowledge to make strides in computational efficiency while retaining their functional correctness.
+This implies that larger models possess a greater capacity to assimilate the nuanced knowledge to make strides in code efficiency while retaining their functional correctness.
 4.5 Failure Analysis
-Table 4: The distribution of failure cases across Code Generation, Code Execution, and Test Case errors. E/M/H indicates Easy/Medium/Hard levels, respectively.
+Table 4 provides an error breakdown of where Code LLMs misstep during the Mercury evaluation:
+(1) Generation Errors arise from syntactical issues.
+The common manifestations include improper indentation, mismatched parentheses, or unexpected truncation.
+Fine-tuning introduces additional knowledge for Code LLMs to adapt the Mercury convention, emphasizing standard indentation, concise code, and minimal comments.
+Therefore, both SFT and DPO generally reduced these errors.
+(2) Execution Errors differ from Generation Errors because they occur after the code has been successfully loaded.
+These errors emerge as exceptions, which could stem from various issues, such as flawed code logic, execution timeouts, memory leakage, or sandbox interruption.
+We observe that SFT tends to aggravate these errors on most models, whereas DPO mitigates these errors successfully.
+(3) Test Case Errors are the most prevalent errors where the code is executed without exceptions, but the output fails to align with the expectation.
+Table 4: The distribution of failure cases across Code Generation, Code Execution, and Test Case errors.
+E/M/H indicates Easy/Medium/Hard levels, respectively.
+We sample 5 solutions for each task, so there are $256*5=1280$ solutions in total for each model.
 Model Name Code Generation Code Execution Test Case Pass
@@
 + DPO 4 12 10 26 76 30 41 40 120 369 277 275
-Table 4 provides an error breakdown of where Code-LLMs misstep during the Mercury evaluation:
-Generation Errors arise from syntactical issues.
-The common manifestations include improper indentation, mismatched parentheses, or unexpected truncation.
-Fine-tuning introduces additional knowledge for Code-LLMs to adapt the Mercury convention, emphasizing standard indentation, concise code, and minimal comments.
-Therefore, both SFT and DPO generally reduced these errors.
-Execution Errors differ from Generation Errors because they occur after the code has been successfully loaded.
-These errors emerge as exceptions, which could stem from various issues, such as flawed code logic, execution timeouts, memory leakage, or sandbox interruption.
-We observe that SFT tends to aggravate these errors on most models, whereas DPO mitigates these errors successfully.
-Test Case Errors are the most prevalent errors where the code is executed without exceptions but the output fails to align with the expectation.
 5 Related Work
@@
 As a pioneer, CodeBLEU [37] adapts the BLEU [35] metric into code generation.
-However, given the abstract nature of programming languages, distinct code can express the equivalent semantics, prompting subsequent benchmarks to harness test cases instead of the similarity measurement.
+However, given the abstract nature of programming languages, distinct code can express the equivalent semantics, prompting subsequent benchmarks to harness test case fuzzing instead of the similarity measurement.
 For example, HumanEval [8] and MBPP [3] consist of hand-written Python programming tasks and corresponding test cases.
@@
 6 Limitations
-In this work, we measure computational efficiency under the assumption that the code runtime is uniformly distributed.
-The simplification allows computational efficiency evaluation via limited solution samples.
+In this work, we measure code efficiency under the assumption that the code runtime is uniformly distributed.
+The simplification streamlines code efficiency evaluation via limited solution samples.
 However, the distribution of code runtime in real-world scenarios is more intricate, which may call for more solution samples to support more precise modeling.
@@
 7 Conclusion
-In this work, we introduced Mercury, the first computational efficiency benchmark for NL2Code evaluation.
-Unlike prior work that focused on functional correctness, our benchmark highlights the importance of computational efficiency.
-By crafting dedicated test case generators and sampling ground-truth solutions across different difficulty levels from Leetcode, we have developed a comprehensive and rigorous Code-LLM evaluation frame.
-We evaluated leading Code-LLMs against benchmarks and found that even though these models are proficient in generating functionally correct code, there is still considerable space for computational efficiency improvement.
-As Code-LLMs become more widely used, computational efficiency determines factual productivity, where Mercury can gauge the vital metric.
+In this work, we introduced Mercury, the first code efficiency benchmark for NL2Code evaluation.
+Unlike prior work that focused on functional correctness, our benchmark highlights the importance of code efficiency.
+By crafting dedicated test case generators and sampling ground-truth solutions across all difficulty levels from Leetcode, we have developed a comprehensive and rigorous Code LLM evaluation frame.
+We evaluated leading Code LLMs against benchmarks and found that even though these models are proficient in generating functionally correct code, there is still considerable space for code efficiency improvement.
+As Code LLMs become more widely used, code efficiency determines factual productivity, where Mercury can gauge the vital metric.
 As a commitment to ongoing research and to foster further innovation in this area, we have open-sourced the Mercury dataset collection framework, laying the groundwork for future advancements in the field.
+Checklist
+  1. 1.
+    For all authors…
+    1. (a)
+      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+[Yes] The abstract and introduction include the main contributions and the research scope.
+    2. (b)
+      Did you describe the limitations of your work? [Yes] See Section 6.
+    3. (c)
+      Did you discuss any potential negative societal impacts of your work?
+[N/A] Given that our work is largely technical and does not engage with societal systems directly, it is unlikely to have negative societal repercussions.
+    4. (d)
+      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] The paper conforms to the ethics review guidelines.
+  2. 2.
+    If you are including theoretical results…
+    1. (a)
+      Did you state the full set of assumptions of all theoretical results? [N/A] The paper does not include any theoretical results.
+    2. (b)
+      Did you include complete proofs of all theoretical results? [N/A] The paper does not include any theoretical results.
+  3. 3.
+    If you ran experiments (e.g. for benchmarks)…
+    1. (a)
+      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
+[Yes] We release our dataset on HuggingFace and our code on GitHub.
+See the abstract footnote.
+    2. (b)
+      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 4.3.
+    3. (c)
+      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Section 4.
+    4. (d)
+      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 4.3 and Appendix Section A.11.
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
+    1. (a)
+      If your work uses existing assets, did you cite the creators? [Yes] See References.
+    2. (b)
+      Did you mention the license of the assets? [Yes] See Section A.13
+    3. (c)
+      Did you include any new assets either in the supplemental material or as a URL? [Yes] See the abstract footnote.
+    4. (d)
+      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Section A.13
+    5. (e)
+      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Section A.13
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects…
+    1. (a)
+      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
+    2. (b)
+      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
+    3. (c)
+      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
 Appendix A Appendix
-A.1 Definition of Mercury Fields
+A.1 Dataset Nutrition Labels
 Table 5: Definitions of the fields within the Mercury dataset.
@@
 Figure 4: Mercury supports two customized data structures: TreeNode and ListNode.
-Table 6: Mercury-eval encompasses 256 tasks, the difficulty of which has been balanced for model evaluation. Mercury-train comprises the remaining 1,633 tasks for model training.
 Splits Easy Medium Hard Sum
@@
 Mercury-eval 88 81 87 256
+Table 6: Mercury-eval encompasses 256 tasks, the difficulty level of which has been balanced for model evaluation. Mercury-train comprises the remaining 1,633 tasks for model training.
 A.3 Sandbox Details
 Time and Memory Limitation.
-Each executed code within the sandbox is subject to certain constraints to ensure fair utilization of resources and to prevent any single program from monopolizing the system.
+Each executed code within the sandbox is subject to certain constraints to ensure fair utilization of resources and to prevent any single code from monopolizing the system resource.
 Specifically, there are two primary constraints: a time limit and a memory limit.
@@
 The memory limit caps the amount of RAM that a process can consume.
-This measure prevents a single program from exhausting the server’s memory resources, which could lead to a denial of service for other users or processes sharing the same infrastructure.
+This measure precludes a single code from exhausting the memory resources, which could lead to a denial of service for subsequent codes.
+In our experiment settings, the timeout limit is 30 seconds, and the memory limit is 2048 MB for each solution execution.
 IO Restriction.
-To mitigate harmful activities such as unauthorized data access or data exfiltration, the sandbox imposes strict input/output (IO) restrictions.
-These restrictions can include limitations on the size and type of files that can be read from or written to the disk and restrictions on the use of network sockets for sending or receiving data.
-By controlling the IO operations, the sandbox can prevent many common exploits and ensure that the code running within it cannot interfere with other processes or the host system.
+To mitigate harmful activities such as unauthorized command execution or data exfiltration, the sandbox imposes strict Input/Output (IO) restrictions.
+These restrictions include limitations on reading from or writing to the disk and restrictions on the use of network sockets for sending or receiving data.
+By controlling the IO operations, the sandbox can prevent many common vulnerabilities and ensure that the code runs without interfering with other processes of the host system.
 Isolated File System.
@@
 The isolated nature of this file system ensures that even if a process within the sandbox attempts to modify or delete files, these changes will not affect the host system or other sandboxes.
-It acts as a layer of security, protecting the host from potential threats and maintaining the integrity of the overall system.
+It acts as a security layer, protecting the host from potential threats and maintaining the integrity of the overall system.
 System Libraries Redirection.
@@
 Single-threaded Evaluation.
-In the context of the sandbox environment, single-threaded evaluation is the process of executing code using only one thread of execution.
-This approach simplifies resource management and timing measurements, avoiding the complexities associated with multi-threaded execution, such as synchronization, race conditions, and deadlocks.
-Single-threading can be particularly important in educational or testing settings where reproducibility and fairness are critical, ensuring that every student or user’s code is evaluated with the same computational resources.
+Single-threaded evaluation refers to executing code using a sole thread of execution, thereby simplifying resource management and timing assessments, and mitigating the intricacies linked with multi-threaded execution, such as synchronization issues, race conditions, and potential deadlocks.
+This mode of operation is especially important in testing environments where reproducibility and fairness are paramount, ensuring that each piece of code is evaluated using identical computational resources.
 Code Efficiency Measurement.
@@
 The final stage involves the sandbox dutifully clearing the namespace environment and the temporary directory.
-Mercury records the time consumed during the stage of Solution instantiation and Test Ease Evaluation as the primary metric for assessing computational efficiency.
+Mercury records the time consumed during the stage of Solution instantiation and Test Ease Evaluation as the primary metric for assessing code efficiency.
 A.4 DPO Experiment Details
@@
 The upper block presents the problem statement with its example, while the subsequent portion exhibits the corresponding solutions.
-Although all solutions provided are functionally correct, they exhibit significant differences in runtimes.
-A.8 Hardware-agnostic Evaluation
-Figure 8: $Beyond$ scores of ‘deepseek-coder-33b’ (solid line) and ‘deepseek-coder-6.7b’ (dashed line) across varied Intel Skylake CPU configurations.
-The results show that $Beyond$ can keep consistent across different hardware configurations.
+Although all solutions are functionally correct, they exhibit significant differences in runtimes.
+A.8 A HumanEval Example
+Figure 8: An HumanEval example of insufficient test cases.
+Even though the code passed all test cases in the dashed-line box, it remains vulnerable to timeout or stack overflow when subjected to a larger input.
 A.9 Prompts for Code Generation
-To guarantee a fair comparison, we apply a unified one-shot prompt template for each pre-trained Code-LLM.
+To guarantee a fair comparison, we apply a unified one-shot prompt template for each pre-trained Code LLM.
 As displayed in Figure 9, the prompt template contains one shot example as well as three placeholders: <task_content>, <code_starter>, and <code_completion>.
@@
 In Mercury experiments, we feed the pretty_content field to the placeholder <task_content>, the prompt field to the placeholder <code_starter>, and the solution field to the placeholder <code_completion>
-NeurIPS Paper Checklist
-The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
-Do not remove the checklist: The papers not including the checklist will be desk rejected.
-The checklist should follow the references and precede the (optional) supplemental material.
-The checklist does NOT count towards the page limit.
-Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
-  • You should answer [Yes] , [No] , or [N/A] .
-  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
-  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
-The checklist answers are an integral part of your paper submission.
-They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
-You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
-The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
-While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
-In general, answering "[No] " or "[N/A] " is not grounds for rejection.
-While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
-All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
-If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
-IMPORTANT, please:
-  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
-  • Keep the checklist subsection headings, questions/answers and guidelines below.
-  • Do not modify the questions and only use the provided macros for your answers.
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [TODO]
-[…]
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [TODO]
-[…]
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [TODO]
-[…]
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [TODO]
-[…]
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [TODO]
-[…]
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [TODO]
-[…]
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [TODO]
-[…]
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [TODO]
-[…]
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [TODO]
-[…]
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [TODO]
-[…]
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [TODO]
-[…]
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [TODO]
-[…]
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [TODO]
-[…]
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [TODO]
-[…]
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [TODO]
-[…]
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+A.10 Hardware-agnostic Evaluation
+Figure 10: Beyond scores of ‘deepseek-coder-33b’ (solid line) and ‘deepseek-coder-6.7b’ (dashed line) across varied Intel Skylake CPU configurations.
+The results show that Beyond can remain consistent across different hardware configurations.
+A.11 Distribution of Bootstrapped Beyond Scores
+Figure 11: Bootstrapped Beyond Distribution.
+We evaluate 3B, 7B, and 15B Starcoder2[24] models using the Mercury benchmark.
+Each model was executed 50 times to ensure score robustness.
+The y-axis in the resulting histogram represents the frequency of observations within each bin.
+A.12 Dataset Metadata
+The Mercury dataset is hosted on Huggingface: https://huggingface.co/datasets/Elfsong/Mercury. The Croissant Metadata can be found at https://huggingface.co/api/datasets/Elfsong/Mercury/croissant.
+A.13 Legal Compliance
+In this study, we have curated a comprehensive dataset by gathering publicly accessible task descriptions and archived solutions from LeetCode (https://leetcode.com/problemset/).
+We have ensured that our collection process is strictly limited to tasks available in the free domain, intentionally excluding any content that falls under the paid services of the platform.
+We abide by Fair Use [33] (Section 107): “the fair use of a copyrighted work, including such use by … scholarship, or research, is not an infringement of copyright”, where fair use is determined by “the purpose and character of the use, including whether such use is of a commercial nature or is for nonprofit educational purposes”.
+With the Mercury dataset, we emphasize its strictly non-commercial nature and underscore its purpose: to facilitate and advance academic research.
+The Mercury dataset is released under Creative Commons Attribution Non Commercial 4.0 [10].